Suppose that a 4 × 4 matrix A has eigenvalues λ1 = 1, λ2 = − 4, λ3 = 6, and λ4 = − 6. Use the following method to find det (A).

If

p(λ) = det (λI − A) = λn + c1λn − 1 + ⋯ + cn

So, on setting λ = 0, we obtain that

det (− A) = cn or det (A) = (− 1)ncn

det (A) = ?


1
Expert's answer
2021-02-24T07:56:47-0500

"Solution: ~\lambda_1=1,~\lambda_2=-4,~\lambda_3=6,~\lambda_4=-6
\\det(\lambda I-A)=\lambda^n +c_1\lambda^{n-1}+c_2 \lambda^{n-2}+....+c_n
\\Here ~ n=4,
\\ \therefore ~det(\lambda I-A)=\lambda^4 +c_1\lambda^3+c_2 \lambda^2+c_3 \lambda+c_4
\\ For \lambda=1,
\\det( I-A)=(1)^4 +c_1(1)^3+c_2 (1)^2+c_3 (1)+c_4
\\~~~~~~~~~~~~~~~~~~~=1 +c_1+c_2 +c_3 +c_4.........................................(I) 
\\ For \lambda=-4,
\\det( -4I-A)=(-4)^4 +c_1(-4)^3+c_2 (-4)^2+c_3 (-4)+c_4
\\~~~~~~~~~~~~~~~~~~~=256 -64c_1+16c_2 -4c_3 +c_4.........................................(II) 
\\ For \lambda=6,
\\det( 6I-A)=(6)^4 +c_1(6)^3+c_2 (6)^2+c_3 (6)+c_4
\\~~~~~~~~~~~~~~~~~~~=1296 +216c_1+36c_2 +6c_3 +c_4.........................................(III)
\\ \\ For \lambda=-6,
\\det( -6I-A)=(-6)^4 +c_1(-6)^3+c_2 (-6)^2+c_3 (-6)+c_4
\\~~~~~~~~~~~~~~~~~~~~~~~~~=1296 -216c_1+36c_2 -6c_3 +c_4.........................................(IV)
\\Re-write~ the ~equation ~(I),(II),(III)~and~(IV),
\\c_4+c_1+c_2 +c_3 =-1..........................................................(I)^*
\\c_4-64c_1+16c_2-4c_3=-256..............................................(II)^*
\\c_4+216c_1+36c_2+6c_3=-1296...........................................(III)^*
\\c_4-216c_1+36c_2-6c_3=-1296...........................................(IV)^*
\\Solving~ (I)^*,(II)^*,(III)^*~and~(IV)^*, we~get~,
\\c_4=144,c_1=3,c_2=-40~and~c_3=-108
\\ \therefore det(\lambda I-A) ~becomes,
\\det(\lambda I-A)=\lambda^4 +3\lambda^3-40 \lambda^2-108 \lambda+144
\\For~ \lambda=0
\\ det(-A)=c_4=144 ~or~det(A)=(-1)^4 .144=144
\\ \therefore det(A)=144
\\Note: Their ~are~ several~ methods ~to ~solve~ system ~of ~linear ~equations."